OpenStudy (anonymous):

What is the factored form of 4k(2) + 20k + 25?

OpenStudy (anonymous):

choices

OpenStudy (anonymous):

4k^2 right ?

OpenStudy (anonymous):

2(k + 5)2 (2k – 5)2 (2k + 5)2 4(k + 5)(k – 5)

OpenStudy (anonymous):

D

OpenStudy (anonymous):

is that 8k in the beginning?

OpenStudy (anonymous):

Yea should it be 8k^2 ?

OpenStudy (anonymous):

oh it's squared I thought it was 4*2

OpenStudy (anonymous):

Nope it's actually A

OpenStudy (anonymous):

Let's factor A and find out ..

OpenStudy (anonymous):

2(k+5)^2 right ?

OpenStudy (anonymous):

I say (2k+5)^2

OpenStudy (anonymous):

great minds think alike

OpenStudy (anonymous):

No your's is actually different though

OpenStudy (anonymous):

hmmmmm Maybe my mind is not so great. And my eyes

OpenStudy (anonymous):

2(k+5)^2 = 2(k^2+10k +25) = 4k^2 + 40k + 100)

OpenStudy (anonymous):

think I'm right on this one

OpenStudy (anonymous):

that's not the correct factored form that the question asks though it's 4k^2 + 20k + 25 so I disagree again

OpenStudy (anonymous):

I used your answer

OpenStudy (anonymous):

slow down and re read your method is correct though

OpenStudy (anonymous):

I expanded your answer

OpenStudy (anonymous):

Okay @norbet123 needs to be more specific with the question and even if you did it your way it's still wrong

OpenStudy (anonymous):

(2k+5)^2

OpenStudy (anonymous):

but if you expand (2k+5)^2, you get the 4k^2+ 10k +25

OpenStudy (anonymous):

which is the original polynomial

OpenStudy (anonymous):

sorry 20k

OpenStudy (anonymous):

Okay I see now sorry I agree now the parenthesis is very important here 2 is inside not outside .. sorry

OpenStudy (anonymous):

All good. :)

OpenStudy (anonymous):

Good Job ;)

OpenStudy (anonymous):

you too. Teamwork

OpenStudy (anonymous):

lol true